Formlas, missing J, L, M, N A. Einstein mass/energy:

E_{0} = mc^{2}

B. Maxwell’s Equations in Vacuum:a

\begin{matrix}∇ · B = 0 \\ ∇ × B = \frac{1}{c} \frac{∂E}{∂t} \end{matrix} \begin{matrix}∇ · E = 0 \\ ∇ × E = −\frac{1}{c} \frac{∂B}{∂t} \end{matrix}

C. Stokes’ Theorem

∫_{M} dω = ∫_{∂M} ω

D.

{Q, Q} = P

E. Primes and the zeta function:

∏_{p} \frac{1}{1 − \frac{1}{p^{s}}} = ∑_{n=1}^{∞} \frac{1}{n^{s}}

F. Heisenberg Uncertainty Principle:

ΔxΔp ≥ ℏ/2

G. Kepler’s Second Law

\frac{dθ}{dt} ∝ \frac{1}{r}

H. Kepler’s Third Law

T ∝ a^{\frac{2}{3}} I. Newton’s Law of Gravitation:

F = \frac{Gm_{1}m_{2}}{r^{2}}

K. Einstein’s General Relativity Equation:

R_{μν} − \frac{1}{2}Rg_{μν} = 8πT_{μν}

O. Yang-Mills equations:

F = dA + A ∧ A
D_{A}^{∗}F_{A} = 0

Non-curved medallion formulas: IV Equations for Lorenz attractor:

\frac{dx}{dt} = σ(y − x)

\frac{dy}{dt} = x(ρ − z) − y

\frac{dz}{dt} = xy − βz

XII. Golden mean = partial fraction expansion:

lim_{n→∞} \frac{F_{n+1}}{F_{n}} = 1 + \frac{1}{1 + …}